# What Did I Do Wrong? Quantifying LLMs’ Sensitivity and Consistency to Prompt Engineering

### v2

@@ Abstract
Federico Errica  and Giuseppe Siracusano  and Davide Sanvito  and Roberto Bifulco NEC Laboratories Europe name.surname@neclab.eu
@@ 1 Introduction
LLMs (Brown et al., 2020) have significantly changed how we process text by providing a straightforward interface, i.e., natural language, to define the problem to be solved (Devlin et al., 2019). They provide software engineers with useful coding tips and can be used as part of larger and more complex software systems (Dakhel et al., 2023). It is common to set up an LLM using a set of instructions called “prompt”, and it soon became clear to both researchers and developers that the prompt itself can greatly influence an LLM’s performance (Zhao et al., 2021; Sclar et al., 2024). The process of writing a good prompt for the current task is called prompt engineering, and a great deal of different techniques have been proposed in this direction (Nori et al., 2023; Sahoo et al., 2024), ranging from a simple description of the problem to providing few-shot examples.
@@
We empirically test these metrics, finding that they indeed convey different information about the LLMs’ behavior. We use this information to qualitatively analyze some of the datasets, mimicking the kind of study that an LLM engineer would do when debugging a system. We also show how these metrics help to improve prompts by identifying different sets of problematic samples.
@@ 2 Related Work
It is well-known that variations of the prompt affect stability of the LLM’s accuracy. In a broad study across diverse tasks, McCoy et al. (2023) showed that LLM performances depend on the likelihood of the input prompt and of the correct output answer. Such result is consistent with LLMs’ autoregressive nature, that is, models trained to maximize a likelihood objective. In light of these considerations, it is not surprising that changing the ordering of the examples in a few-shot prompting strategy can lead to almost random accuracy on sentiment analysis tasks (Zhao et al., 2021), or that the nature of prompts affects LLMs’ benchmarks (Ailem et al., 2024). Similar considerations motivated frameworks like FormatSpread, which predicts the expected performance under prompt’s variations without accessing the LLM’s weights (Sclar et al., 2024). Also, spurious features in the prompt have severe repercussions on security vulnerabilities’ detection, where LLMs appear inconsistent and unfaithful (Ullah et al., 2024), while Yang et al. (2024) study the effect of prompt rephrasings and LLMs’ temperature on classification and uncertainty metrics. Finally, we mention the comprehensive benchmark of Liang et al. (2023), where metrics such as invariance to semantic-preserving transformations are computed w.r.t. accuracy changes. Note that the solution to the spurious feature problem might also reside in a more structured approach: Retrieval Augmented Generation (Lewis et al., 2020) or Knowledge Graphs-enhanced LLMs (Luo et al., 2024) reduce hallucinations and prompt dependencies.
@@
It does not fall within the scope of this paper to analyze output uncertainty. Because LLMs can also behave pseudo-deterministically (Yang et al., 2024), this limits the impact of some probability-based uncertainty quantification metrics. For a comprehensive study of uncertainty quantification techniques on LLMs, such as verbalization and generation likelihood, we refer the reader to Huang et al. (2023) and Geng et al. (2024).
@@ 3 Methodology
In this work, the sampler $\mathcal{S}$ is an LLM that is tasked to rephrase prompts (Appendix D shows examples), but it could also be an algorithm that generates text modifications as in Liang et al. (2023). We then define the sensitivity of an LLM to the prompt to reflect how much the LLM prediction varies under the rephrasings of the original prompt.
@@ 4 Experiments
The goal of the experiments is to assess how sensitivity and consistency can help in the context of English-based text classification datasets, analyzing the impact of prompt variations and prompting strategies from a different angle compared to previous works, but most importantly showing developers how to use these metrics [^2] The code to reproduce our results is provided in the supplementary material and will be open-sourced at publication.. Therefore, it is outside the scope of this work to perform an extensive benchmark of these metrics across many datasets and LLMs, which would not be helpful to our purpose. large number of LLMs across We ran the LLMs on a 4 Tesla V100 server with 32 GBs of memory and 252 GBs of RAM.
@@
We consider three different prompting strategies: a simple strategy, where the prompt consists of the task description and the list of classes; a detail strategy, where we provide a detailed description of each class; and a 1-shot strategy in which, compared to simple, we also provide one example for each class taken from samples that do not belong to the test set. To build different rephrasings of the task description only, we use the aforementioned LLMs. In particular, the prompt asks to rephrase the task description by changing the length or adding unnecessary words as long as the meaning remains the same. For the purpose of our work, which is only to show how sensitivity and consistency can be used, we fix $Q=30$ after observing that increasing this value did not vary results significantly and was enough for our demonstrative purposes. Also, the “right” value of $Q$ ultimately depends on the user’s compromise between computational resources and stability of resulting statistics. In the quantitative analyses, we display the average sensitivity and average consistency, as well as the micro F1-score across all $\rho$ for completeness. Since the distributions of values are far from being Gaussian, we do not report standard deviations and refer to more qualitative analyses in later sections that show deviations from the mean values.
@@ 6 Conclusions
In this work, we have expanded the LLM developer’s toolbox with two diagnostic metrics, namely sensitivity and consistency, that complement performance metrics such as accuracy. In our experiment, we showed how different prompting strategies influence these metrics and how we can guide prompt engineering decisions based on new criteria that value intrinsic characteristics of LLMs’ predictions. Indeed, an LLM that is very sensitive to variation of the prompt and has high test set accuracy might not be a good choice in a production environment, where multiple intermediate steps exist and each could lead to minor alterations in the prompt of the LLM predictor. Notably, sensitivity does not require access to the ground truth labels, and it would be interesting in future work to extend it to tasks different from classification, for instance, code generation. Also, while this work mentioned prompt optimization, it is still an open question how to integrate these metrics in an automatic prompt engineering framework, leading to LLMs insensitive to nonsensical prompt variations and consistent in their (good) performances. As LLM will continue to be neural autoregressive models, and more generally pure neural networks, it is likely that the sensitivity to the input will remain a pain point of these architectures.
@@ Appendix C Using Consistency to Improve Prompts
Appendix D Examples of Prompt Rephrasings
For completeness, we report below some task rephrasings that we generated for the different datasets using LLama3. The complete list of rephrasings for the other models is similar and can be inspected in the supplementary material. Please note that these sentences are just a part of the whole prompt and they are used to describe the task; the rest of the prompt remains the same for all models. We did not observe strange behavior in the rephrasings produced by the LLM. The first sentence corresponds to the original prompt.

### v3

@@ Abstract
Federico Errica  and Giuseppe Siracusano  and Davide Sanvito  and Roberto Bifulco NEC Italia  and NEC Laboratories Europe name.surname@neclab.eu
@@ 1 Introduction
LLMs (Brown et al., 2020) have significantly changed how we process text by providing a straightforward interface, i.e., natural language, to define the problem to be solved (Devlin et al., 2019). They provide software engineers with useful coding tips and can be used as part of larger and more complex software systems (Dakhel et al., 2023). It is common to set up an LLM using a set of instructions called “prompt”, and it soon became clear to both researchers and developers that the prompt itself can greatly influence an LLM’s performance (Zhao et al., 2021; Sclar et al., 2024; Yang et al., 2024b). The process of writing a good prompt for the current task is called prompt engineering, and a great deal of different techniques have been proposed in this direction (Nori et al., 2023; Sahoo et al., 2024), ranging from a simple description of the problem to few-shot examples.
@@
We empirically test these metrics, finding that they indeed convey different information about the LLMs’ behavior. We use this information to qualitatively analyze some of the datasets, mimicking what an LLM engineer would do when debugging a real system. We also show how these metrics help to improve prompts by identifying different sets of problematic samples.
@@ 2 Related Work
It is well-known that variations of the prompt affect stability of the LLM’s accuracy. In a broad study across diverse tasks, McCoy et al. (2023) showed that LLM performances depend on the likelihood of the input prompt and of the correct output answer. Such result is consistent with LLMs’ autoregressive nature, that is, models trained to maximize a likelihood objective. In light of these considerations, it is not surprising that changing the ordering of the examples in a few-shot prompting strategy can lead to almost random accuracy on sentiment analysis tasks (Zhao et al., 2021), or that the nature of prompts affects LLMs’ benchmarks (Ailem et al., 2024). Similar considerations motivated frameworks like FormatSpread, which predicts the expected performance under prompt’s variations without accessing the LLM’s weights (Sclar et al., 2024). Also, spurious features in the prompt have severe repercussions on security vulnerabilities’ detection, where LLMs appear inconsistent and unfaithful (Ullah et al., 2024), while Yang et al. (2024a) study the effect of prompt rephrasings and LLMs’ temperature on classification and uncertainty metrics. Finally, we mention the comprehensive benchmark of Liang et al. (2023), where metrics such as invariance to semantic-preserving transformations are computed w.r.t. accuracy changes. Note that the solution to the spurious feature problem might also reside in a more structured approach: Retrieval Augmented Generation (Lewis et al., 2020) or Knowledge Graphs-enhanced LLMs (Luo et al., 2024) reduce hallucinations and prompt dependencies.
@@
It does not fall within the scope of this paper to analyze output uncertainty. Because LLMs can also behave pseudo-deterministically (Yang et al., 2024a), this limits the impact of some probability-based uncertainty quantification metrics. For a comprehensive study of uncertainty quantification techniques on LLMs, such as verbalization and generation likelihood, we refer the reader to Huang et al. (2023) and Geng et al. (2024).
@@ 3 Methodology
In this work, the sampler $\mathcal{S}$ is an LLM that is tasked to rephrase prompts (Appendix E shows examples), but it could also be an algorithm that generates text modifications as in Liang et al. (2023). We then define the sensitivity of an LLM to the prompt to reflect how much the LLM prediction varies under the rephrasings of the original prompt.
@@ 4 Experiments
The goal of the experiments is to assess how sensitivity and consistency can help in the context of English-based text classification datasets, analyzing the impact of prompt variations and prompting strategies from a different angle compared to previous works, but most importantly showing developers how to use these metrics [^2] The code is available at: https://github.com/nec-research/sensitivity-consistency-LLM.. Therefore, it is outside the scope of this work to perform an extensive benchmark of these metrics across many datasets and LLMs, which would not be helpful to our purpose. large number of LLMs across We ran the LLMs on a 4 Tesla V100 server with 32 GBs of memory and 252 GBs of RAM.
@@
We analyze three different prompting strategies: a simple strategy, where the prompt consists of the task description and the list of classes; a detail strategy, where we provide a detailed description of each class; and a 1-shot strategy in which, compared to simple, we also provide one example for each class taken from samples that do not belong to the test set. To build different rephrasings of the task description only, we use the aforementioned LLMs. In particular, the prompt asks to rephrase the task description by changing the length or adding unnecessary words as long as the meaning remains the same. For the purpose of our work, which is only to show how sensitivity and consistency can be used, we fix $Q=30$ after observing that increasing this value did not vary results significantly and was enough for our demonstrative purposes. Also, the “right” value of $Q$ ultimately depends on the user’s compromise between computational resources and stability of resulting statistics. In the quantitative analyses, we display the average sensitivity and average consistency, as well as the micro F1-score across all $\rho$. Sensitivity and micro F1-score are averaged across samples, whereas pair-wise consistencies are averaged together. Since the distributions of sensitivity and consistency values are far from being Gaussian, looking at certain statistics such as the standard deviation can convey a misleading message of instability. Therefore, we will perform qualitative analyses in later sections that show large deviations from the mean values, but in the interest of completeness we report standard deviations in Appendix D.
@@ 6 Conclusions
In this work, we have expanded the LLM developer’s toolbox with two diagnostic metrics, namely sensitivity and consistency, that complement performance metrics such as accuracy. In our experiment, we showed how different prompting strategies influence these metrics and how we can guide prompt engineering decisions based on new criteria that value intrinsic characteristics of LLMs’ predictions. Indeed, an LLM that is very sensitive to variation of the prompt and has high test set accuracy might not be a good choice in a production environment, where multiple intermediate steps exist and each could lead to minor alterations in the prompt of the LLM predictor. Notably, sensitivity does not require access to the ground truth labels, and it would be interesting in future work to extend it to tasks different from classification, for instance, code generation. Also, while this work mentioned prompt optimization, it is still an open question how to integrate these metrics in an automatic prompt engineering framework, leading to LLMs insensitive to nonsensical prompt variations and consistent in their (good) performances. Our hope is that sensitivity and consistency to the input will become relevant metrics from both an academic and industrial perspective, helping to identify pain points of LLMs.
@@ Appendix C Using Consistency to Improve Prompts
Appendix D Analysis of Standard Deviation Values
For completeness, we also report standard deviation values of our results. Please notice, as we write in Section 4, that the distribution of results is not Gaussian and there is no reason it should be; therefore, statistics such as the standard deviation can convey misleading information. This is why we decided not to provide these numbers in the main paper and rather focus more on qualitative investigations, that analyze the actual distributions of sensitivity and consistency values and allow for a more meaningful analysis of failure cases.
Llama3 Mixtral
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
TREC .148/.339/.085 .149/.339/.080 .146/.351/.051 .168/.305/.071 .182/.316/.056 .158/.341/.050
CB .082/.163/.005 .073/.156/.004 .067/.170/.006 .216/.351/.048 .188/.367/.049 .188/.360/.076
RTE .244/.397/.066 .289/.389/.121 .160/.402/.013 .277/.376/.037 .263/.383/.041 .152/.401/.010
DBPedia .060/.202/.011 .095/.217/.062 .076/.210/.031 .104/.321/.025 .113/.332/.014 .115/.329/.020
WoS .103/.466/.008 .091/.472/.008 .092/.472/.008 .164/.424/.019 .171/.416/.013 .171/.401/.028
GPT-3.5 GPT-4o
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
Simple
$S_{\tau}$/$C_{y}$/F1
Detail
$S_{\tau}$/$C_{y}$/F1
1-shot
$S_{\tau}$/$C_{y}$/F1
TREC .174/.387/.026 .164/.371/.034 .155/.385/.018 .099/.344/.013 .120/.367/.030 .096/.335/.010
CB .277/.284/.028 .265/.252/.029 .252/.260/.025 .174/.313/.012 .164/.306/.013 .145/.266/.014
RTE .229/.267/.023 .232/.245/.025 .234/.309/.015 .295/.300/.053 .187/.305/.006 .168/.302/.006
DBPedia .071/.230/.007 .081/.258/.012 .070/.228/.008 .069/.201/.015 .068/.229/.015 .057/.230/.007
WoS .140/.430/.018 .138/.431/.018 .141/.421/.035 .105/.461/.004 .100/.463/.003 .100/.463/.003
Table 2: Standard deviation values of sensitivity, consistency, and micro-F1 score across all prompt rephrasings $\rho$ are shown for different datasets, models, and prompting strategies. We remind the reader that standard deviation values can be misleading when the distribution is not Gaussian (which is the case in our experiments).
Appendix E Examples of Prompt Rephrasings
We report below some task rephrasings that we generated for the different datasets using LLama3. The complete list of rephrasings for the other models is similar and can be inspected in the supplementary material. Please note that these sentences are just a part of the whole prompt and they are used to describe the task; the rest of the prompt remains the same for all models. We did not observe strange behavior in the rephrasings produced by the LLM. The first sentence corresponds to the original prompt.